Question 2 1 pts A 1,500,000 SF plant employs 300 people and produces 203,000 LCD panel per month. The plant is idle, on average, 2 months each year for maintenance, repairs, and retooling. The plant has fixed costs of $4,000,000 per month and variable costs of $70 per LCD panel. What is the total operating cost for the plant per YEAR? Express your answer in $ to the nearest $1 million
Answer #1

Annual fixed cost = 4000000*12 = $48000000

Monthly variable cost = 203000*70 = $14210000

No. of months of production in a year= 10

annual variable cost = 14210000*10 = $142100000

So,

Annual operating cost = 48000000+142100000

Annual operating cost = $190.1 Million or $190 Million